A test consists of 10​ true/false questions. To pass the test a student must answer at least 6 questions correctly. If a student guesses on each​ question, what is the probability that the student will pass the​ test? Round to three decimal places.

A.

0.377

B.

0.828

C.

0.205

D.

0.172
Answer #1

We have here given,

n=10 and p=1/2=0.5

=1-0.6230....................by using Excel command =BINOMDIST(5,10,0.5,TRUE)

=0.377

Option A) is correct.